MAKEFILE TIPS (VERSION 0.0.2)

◆1. はじめに (copy right など)

この文書はもともと (主として GNU 版の) make について、使い方を自分用の忘備録としてまとめたものです。この文書は書きかけです。間違った事が書いてあって、この通りに行うと大事なファイルを失う可能性が十分あります。読んでいる皆さんはそのあたりを御理解の上、御自身でよく実験してからお使い下さることをお勧めします。また、この文書をよむには (いまのところ) make などの知識がある程度必要です。

この文書は GNU Free Documentation License (version 1.1 あるいはそれ以降) に従います。

   (c) Copyright 2000 Yoshifumi Tsuchimoto
    Permission is granted to copy, distribute and/or modify this document
    under the terms of the GNU Free Documentation License, Version 1.1
    or any later version published by the Free Software Foundation;
    with no Invariant Sections, with no  Front-Cover Texts,
    and with no Back-Cover Texts.
    A copy of the license is included in the section entitled "GNU
    Free Documentation License".

◆2. 意外に便利な make

make はプログラミングの際には不可欠な道具です。 Makefile に動作を記述しておくと、 make コマンドを実行するだけで様々なことを自動実行してくれます。

プログラミングに限らず、同じことを何度も繰り返すような状況ではこの自動実行はとても有用で、実は私 (つちもと) は platex で数学セミナーの原稿を書く時に make に大変お世話になりました。具体的には、次のような目的に make を利用しました。

1. 不要な .bak ファイルなどを消す。

2. mupad で生成した画像を別形式に直す。

3. ディレクトリ全体のスナップショットをとる。

make はこれらの操作で実際に手を下している「何でも屋」であるというわけではありません。そうではなくて、これらを行うためのいろいろなツールを呼び出して指図する「指揮者」として振舞うのです。

よいオーケストラなら指揮者がいなくても一応は演奏ができるように、 make なしでも上記操作を行えますが、やはり指揮者である make があった方が効率面で優ります。

しかも Makefile に「操作法」を書いておくことで、「操作法を忘れる」危険性をかなり減らすことができるのです。

そのときの Makefile (一部省略してある) は最後の方の節に載せた Makefile です。この tips ではこの Makefile をつかったときの経験の幾つかをお話しする事にしましょう。 ◆3. 一般的な注意

1. Makefile を書くときにはタブが重要な意味を持ちます。不注意に消してしまったり、空白で置き換えてしまうと全く動かなくなってしまいます。カットアンドペーストを行う際にはとくに注意が必要でしょう。

2. make はカレントディレクトリの Makefile を読み込みます。 (ほかにも makefile や,(GNU make なら) GNUmakefile などを既定の順番で読み込みます。詳細についてはマニュアルを参照してみてください。 ) 見知らぬディレクトリ (や、他人に makefile 等を勝手に書かれたり、書き換えられたりしてしまう可能性のある環境) で make コマンドを実行するのはセキュリティー的には自殺行為に等しいものです。くれぐれもご注意下さい。

◆4. いろいろなターゲット

make を実行したときにやってもらいたい「目標」のことをターゲットと呼びます。ここでは比較的よく用いられるターゲットについてみていきましょう。

(なお、 clean や all などの名前は慣用になってはいますが、絶対これでなければならないと言うわけでもありません。 )

4.1. clean から始めよう

(警告) この節から始めるのは、簡単で、しかも有用だからですが、実は危険でもあります。不注意にファイルが消えてしまう事のないように、テスト用のディレクトリで 3 回以上チェックしてから応用するようにして下さい。

エディタの bak ファイルなど、ハードディスクのいらないゴミは知らないうちにどんどんたまっていくものです。そうかといって rm コマンドを乱発していると、必要なファイルまで思わず消してしまったりします。そんなとき、作業しているディレクトリに

all:
clean:
        rm -f *.bak

という内容の Makefile を作っておくと便利です。 (rm の行の先頭は TAB です。 ) make clean と打つだけで .bak ファイルを全て消してくれます。 all: の行も重要です。これがないと単に make と打っただけでも make clean したのと同

じになってしまいます。

.bak ファイルだけではなくて、いろいろなゴミがでて来る場合には、次のような Makefile を使えばよいでしょう。

all:
clean:
        rm -f *.bak
rm -f *~
rm -f core
rm -f \#*\#

たかがゴミを消すだけなのに、わざわざ Makefile を作るのをいぶかる方もいるかも知れませんが、少なくとも次のような利点があるのです。

1. コマンドの打ち間違いの心配が減る。 (rm -f * ~ 等と打ってしまったら悲惨ですから。 )

2. どれがゴミか消す都度悩む必要がない。

もちろん、最初に Makefile を作る際にはどれを消すべきか慎重に判断して、 Makefile の打ち間違いもないように気を配らなければなりません。 (慣れないうちは、消すファイルは極力控えめにして、絶対いらないファイルに限定するようにした方がいいでしょう。 )

1 年に及ぶ原稿書きの場合など、これがあると無しとでは大違いなのが、理解頂けるでしょうか。

(※) なお、一般に流布しているプログラムのソースでは、 make clean すると make で作ったものを洗いざらい消してしまう (プログラムのソースなどの「材料」は消さない) 動作をするようになっています。ここではそれをちょっと違う用途に「流用」しているわけです。

4.2. すべてをつくる (all)

単に make と打ったときは、 Makefile の先頭にあるターゲットを実行します。通常はそのターゲットは all という名前が付けられます。

たとえば Makefile の先頭に

all:
platex a.tex

と書いておけば、 make と打つだけで platex a.tex というコマンドを実行したのと同じ効果が得られるわけです。

これぐらいだったら単にコマンドを打っても変わらない、と感じる方もいらっしゃるかも知れませんが、

1. コマンド名 (platex) やファイル名 (a.tex) を覚えたり、確かめたりする手間が減る。

2. vim の :make コマンドなどでエディタから直接使える。

という利点があります。 (:make コマンドは elvis 等にも実装されているようです。 vi 自身にあったかどうかはちょっと調べていないです。 ) おそらく emacs でも適当な方法で make を呼び出せることと思います。

とにかくいつでもどこでも make で自分の作ったファイルが必要に応じて加工されるわけですから、楽です。おかげで私は「the」より速く「make」とタイプできるようになってしまいました。

4.3. インストールする (install)

clean, all ときて、あと頻繁に目にするターゲットと言えば、 install でしょう。実際、近頃の GNU のソフトウエアの多くは、ソースを展開した後、

./configure
make
make install

でコンパイルからインストールまでしてしまいます。 (最初の ./configure は Makefile を使用しているコンピュータの環境等に合わせてつくるためのシェルスクリプトです。 )

プライベートなものをプライベートなディレクトリに単においておきたい場合には、 install ターゲットはあまり必要ないので、私の Makefile には、 install ターゲットはありません。 (その代わりに、私のコンピュータからサーバにファイルを転送する「rsync」というターゲットを作ったのですが、これは場合に応じて記述が著しく異なる (しかもプライベートなことでもある) ので削除してあります。

4.4. ターゲットとしてのファイル

本当はこちらが王道で、 make の解説書などではこちらから解説してあります。これについては後で述べることにしましょう。 4.5. その他のターゲット

ターゲットは自由に、好きなだけ作ることができます。たとえば、私の Makefile では、 make print で印刷ができ、 make snapshot で現在のディレクトリのスナップショットをとることができるようになっています。

◆5. サフィックスルール

TeX ユーザなら、 .dvi ファイルは .tex ファイルをコンパイルして作ることを知っていますし、 C を使っている人は .c ファイルから .o を作ることをご存知でしょう。このように「 . なんとか」から「. かんとか」を作るルールがサフィックスルールです。

5.1. .SUFFIXES:を準備しよう

まず最初に、サフィックスの「順序関係」を make におしえてやる必要があります。 make はデフォルトで幾つかの順序を知っていますが、これはまあ手っ取り早く一時的な

仕事を片付けるときのもので、本格的なプロジエクトを組みたいとき、いろいろな環境で Makefile を使い回すとき、はたまた私のようにプログラミングとは全く関係ない状況で make を使いたいときなどにはどうしてもこれが必要になります。忘れずに指定しておきましょう。書きかたはこんな感じです

.SUFFIXES:
.SUFFIXES:  .ps  .obj .dvi .png .mupad .tex .sed

一行目はデフォルトのサフィックスの順序を忘れなさい、という指定です。二行目で改めてサフィックスの順序をいれています。右の方にいく程「原材料」に近く、左の方にいく程「製品」に近いと思えばよいでしょう。

◆6. 原稿を打つのにつちもとが使った makefile

# makefile for ' 数学セミナー' tex files
SHELL = /bin/sh
.PHONY: expose
.SUFFIXES:
.SUFFIXES:  .ps  .obj .dvi .png  .mupad .tex .sed
.PRECIOUS: %.obj %.ps
X=$(shell \
count=01; \
cd ..; while [ -f $(TARGET)-$$count.tar.gz ] ;do \
x=`expr $$count + 1`;\
count=`printf "%02d" $$x` ; done; echo $$count)
REL=$(shell \
count=01; \
cd ..; while [ -f $(TARGET)-rel$$count.tar.gz ] ;do \
x=`expr $$count + 1`;\
count=`printf "%02d" $$x` ; done; echo $$count)

TARGET :=xx12
EPSFILES :=fourD.eps fourD2.eps sqrt.eps sqrt2.eps tsunageru.eps \
tangent.eps graph.eps exp0.eps cover.eps
##################################################################
all:: $(TARGET).ps.gz
html:: html_stamp
html_stamp: $(TARGET).ps.gz
/usr/local/latex2html/latex2html -no_math $(TARGET).tex
touch html_stamp
clean:
rm -f *~
rm -f core
rm -f \#*\#
rm -f *.bak
rm -f .??*.bak
rm -f *.dvi_tmp *.eps_tmp
rm -f *.beuc
.tex.dvi:: super.bib
platex $*.tex
zsh -c '! grep -c " Rerun to" $*.log >/dev/null || platex $*.tex '
.mupad.png:
xmupad $*.mupad
.png.obj::
echo 'please run tgif and make $*.obj' && false
.obj.ps::
tgif -print -ps -color  $*.obj
.dvi.ps::
dvips  $*.dvi
%.eps: %.ps %.ptex %.sed
platex $*.ptex; mv -f $*.dvi $*.dvi_tmp
dvips -f -E $*.dvi_tmp >$*.eps_tmp
rm -f $*.dvi_tmp
sed -f $*.sed $*.eps_tmp >$*.eps
rm -f $*.eps_tmp
preview: $(TARGET).dvi
xdvi -s 5 -expert -nopostscript  $(TARGET) 1>/dev/null 2>/dev/null &
psview: $(TARGET).ps
gv -antialias  $(TARGET) 1>/dev/null 2>/dev/null &
expose:
cp -f $(TARGET).tex expose/
cp -f $(TARGET).aux expose/
cp -f $(TARGET).dvi expose/
chmod a+r expose/*.dvi
make -C expose html
print: $(TARGET).dvi
lpr $(TARGET).dvi
links:

scripts/makelinks.sh
$(TARGET).dvi: $(EPSFILES) stddefs.tex
$(TARGET).ps: $(TARGET).dvi
$(TARGET).ps.gz:$(TARGET).ps
gzip -c -9 $(TARGET).ps >$(TARGET).ps.gz  2>/dev/null &
snapshot:
scripts/makelinks.sh
make html
echo $X
cd ..; tar -zcvf $(TARGET)-$X.tar.gz $(TARGET)
echo ' スナップショット$X \
<a href=$(TARGET)-$X.tar.gz>$(TARGET)-$X.tar.gz</a><br>'\
>>../list$(TARGET).html
release:
scripts/makelinks.sh
make html
echo $X
cd ..; tar -zcvf $(TARGET)-rel$(REL).tar.gz $(TARGET)
echo ' リリース候補$(REL) \
<a href=$(TARGET)-rel$(REL).tar.gz>$(TARGET)-rel$(REL).tar.gz</a><br>'\
>>../list$(TARGET).html

◆7. GNU Free Documentation License

GNU Free Documentation License
   Version 1.1, March 2000
 Copyright (C) 2000  Free Software Foundation, Inc.
     59 Temple Place, Suite 330, Boston, MA  02111-1307  USA
 Everyone is permitted to copy and distribute verbatim copies
 of this license document, but changing it is not allowed.

0. PREAMBLE
The purpose of this License is to make a manual, textbook, or other
written document "free" in the sense of freedom: to assure everyone
the effective freedom to copy and redistribute it, with or without
modifying it, either commercially or noncommercially.  Secondarily,
this License preserves for the author and publisher a way to get
credit for their work, while not being considered responsible for
modifications made by others.
This License is a kind of "copyleft", which means that derivative
works of the document must themselves be free in the same sense.  It

complements the GNU General Public License, which is a copyleft
license designed for free software.
We have designed this License in order to use it for manuals for free
software, because free software needs free documentation: a free
program should come with manuals providing the same freedoms that the
software does.  But this License is not limited to software manuals;
it can be used for any textual work, regardless of subject matter or
whether it is published as a printed book.  We recommend this License
principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS
This License applies to any manual or other work that contains a
notice placed by the copyright holder saying it can be distributed
under the terms of this License.  The "Document", below, refers to any
such manual or work.  Any member of the public is a licensee, and is
addressed as "you".
A "Modified Version" of the Document means any work containing the
Document or a portion of it, either copied verbatim, or with
modifications and/or translated into another language.
A "Secondary Section" is a named appendix or a front-matter section of
the Document that deals exclusively with the relationship of the
publishers or authors of the Document to the Document's overall subject
(or to related matters) and contains nothing that could fall directly
within that overall subject.  (For example, if the Document is in part a
textbook of mathematics, a Secondary Section may not explain any
mathematics.)  The relationship could be a matter of historical
connection with the subject or with related matters, or of legal,
commercial, philosophical, ethical or political position regarding
them.
The "Invariant Sections" are certain Secondary Sections whose titles
are designated, as being those of Invariant Sections, in the notice
that says that the Document is released under this License.
The "Cover Texts" are certain short passages of text that are listed,
as Front-Cover Texts or Back-Cover Texts, in the notice that says that
the Document is released under this License.
A "Transparent" copy of the Document means a machine-readable copy,
represented in a format whose specification is available to the
general public, whose contents can be viewed and edited directly and
straightforwardly with generic text editors or (for images composed of

pixels) generic paint programs or (for drawings) some widely available
drawing editor, and that is suitable for input to text formatters or
for automatic translation to a variety of formats suitable for input
to text formatters.  A copy made in an otherwise Transparent file
format whose markup has been designed to thwart or discourage
subsequent modification by readers is not Transparent.  A copy that is
not "Transparent" is called "Opaque".
Examples of suitable formats for Transparent copies include plain
ASCII without markup, Texinfo input format, LaTeX input format, SGML
or XML using a publicly available DTD, and standard-conforming simple
HTML designed for human modification.  Opaque formats include
PostScript, PDF, proprietary formats that can be read and edited only
by proprietary word processors, SGML or XML for which the DTD and/or
processing tools are not generally available, and the
machine-generated HTML produced by some word processors for output
purposes only.
The "Title Page" means, for a printed book, the title page itself,
plus such following pages as are needed to hold, legibly, the material
this License requires to appear in the title page.  For works in
formats which do not have any title page as such, "Title Page" means
the text near the most prominent appearance of the work's title,
preceding the beginning of the body of the text.

2. VERBATIM COPYING
You may copy and distribute the Document in any medium, either
commercially or noncommercially, provided that this License, the
copyright notices, and the license notice saying this License applies
to the Document are reproduced in all copies, and that you add no other
conditions whatsoever to those of this License.  You may not use
technical measures to obstruct or control the reading or further
copying of the copies you make or distribute.  However, you may accept
compensation in exchange for copies.  If you distribute a large enough
number of copies you must also follow the conditions in section 3.
You may also lend copies, under the same conditions stated above, and
you may publicly display copies.

3. COPYING IN QUANTITY
If you publish printed copies of the Document numbering more than 100,
and the Document's license notice requires Cover Texts, you must enclose
the copies in covers that carry, clearly and legibly, all these Cover

Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on
the back cover.  Both covers must also clearly and legibly identify
you as the publisher of these copies.  The front cover must present
the full title with all words of the title equally prominent and
visible.  You may add other material on the covers in addition.
Copying with changes limited to the covers, as long as they preserve
the title of the Document and satisfy these conditions, can be treated
as verbatim copying in other respects.
If the required texts for either cover are too voluminous to fit
legibly, you should put the first ones listed (as many as fit
reasonably) on the actual cover, and continue the rest onto adjacent
pages.
If you publish or distribute Opaque copies of the Document numbering
more than 100, you must either include a machine-readable Transparent
copy along with each Opaque copy, or state in or with each Opaque copy
a publicly-accessible computer-network location containing a complete
Transparent copy of the Document, free of added material, which the
general network-using public has access to download anonymously at no
charge using public-standard network protocols.  If you use the latter
option, you must take reasonably prudent steps, when you begin
distribution of Opaque copies in quantity, to ensure that this
Transparent copy will remain thus accessible at the stated location
until at least one year after the last time you distribute an Opaque
copy (directly or through your agents or retailers) of that edition to
the public.
It is requested, but not required, that you contact the authors of the
Document well before redistributing any large number of copies, to give
them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS
You may copy and distribute a Modified Version of the Document under
the conditions of sections 2 and 3 above, provided that you release
the Modified Version under precisely this License, with the Modified
Version filling the role of the Document, thus licensing distribution
and modification of the Modified Version to whoever possesses a copy
of it.  In addition, you must do these things in the Modified Version:
A. Use in the Title Page (and on the covers, if any) a title distinct
   from that of the Document, and from those of previous versions
   (which should, if there were any, be listed in the History section
   of the Document).  You may use the same title as a previous version
   if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities
responsible for authorship of the modifications in the Modified
Version, together with at least five of the principal authors of the
Document (all of its principal authors, if it has less than five).
C. State on the Title page the name of the publisher of the
Modified Version, as the publisher.
D. Preserve all the copyright notices of the Document.
E. Add an appropriate copyright notice for your modifications
adjacent to the other copyright notices.
F. Include, immediately after the copyright notices, a license notice
giving the public permission to use the Modified Version under the
terms of this License, in the form shown in the Addendum below.
G. Preserve in that license notice the full lists of Invariant Sections
and required Cover Texts given in the Document's license notice.
H. Include an unaltered copy of this License.
I. Preserve the section entitled "History", and its title, and add to
it an item stating at least the title, year, new authors, and
publisher of the Modified Version as given on the Title Page.  If
there is no section entitled "History" in the Document, create one
stating the title, year, authors, and publisher of the Document as
given on its Title Page, then add an item describing the Modified
Version as stated in the previous sentence.
J. Preserve the network location, if any, given in the Document for
public access to a Transparent copy of the Document, and likewise
the network locations given in the Document for previous versions
it was based on.  These may be placed in the "History" section.
You may omit a network location for a work that was published at
least four years before the Document itself, or if the original
publisher of the version it refers to gives permission.
K. In any section entitled "Acknowledgements" or "Dedications",
preserve the section's title, and preserve in the section all the
substance and tone of each of the contributor acknowledgements
and/or dedications given therein.
L. Preserve all the Invariant Sections of the Document,
unaltered in their text and in their titles.  Section numbers
or the equivalent are not considered part of the section titles.
M. Delete any section entitled "Endorsements".  Such a section
may not be included in the Modified Version.
N. Do not retitle any existing section as "Endorsements"
or to conflict in title with any Invariant Section.
If the Modified Version includes new front-matter sections or
appendices that qualify as Secondary Sections and contain no material
copied from the Document, you may at your option designate some or all
of these sections as invariant.  To do this, add their titles to the
list of Invariant Sections in the Modified Version's license notice.
These titles must be distinct from any other section titles.

You may add a section entitled "Endorsements", provided it contains
nothing but endorsements of your Modified Version by various
parties--for example, statements of peer review or that the text has
been approved by an organization as the authoritative definition of a
standard.
You may add a passage of up to five words as a Front-Cover Text, and a
passage of up to 25 words as a Back-Cover Text, to the end of the list
of Cover Texts in the Modified Version.  Only one passage of
Front-Cover Text and one of Back-Cover Text may be added by (or
through arrangements made by) any one entity.  If the Document already
includes a cover text for the same cover, previously added by you or
by arrangement made by the same entity you are acting on behalf of,
you may not add another; but you may replace the old one, on explicit
permission from the previous publisher that added the old one.
The author(s) and publisher(s) of the Document do not by this License
give permission to use their names for publicity for or to assert or
imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS
You may combine the Document with other documents released under this
License, under the terms defined in section 4 above for modified
versions, provided that you include in the combination all of the
Invariant Sections of all of the original documents, unmodified, and
list them all as Invariant Sections of your combined work in its
license notice.
The combined work need only contain one copy of this License, and
multiple identical Invariant Sections may be replaced with a single
copy.  If there are multiple Invariant Sections with the same name but
different contents, make the title of each such section unique by
adding at the end of it, in parentheses, the name of the original
author or publisher of that section if known, or else a unique number.
Make the same adjustment to the section titles in the list of
Invariant Sections in the license notice of the combined work.
In the combination, you must combine any sections entitled "History"
in the various original documents, forming one section entitled
"History"; likewise combine any sections entitled "Acknowledgements",
and any sections entitled "Dedications".  You must delete all sections
entitled "Endorsements."

6. COLLECTIONS OF DOCUMENTS
You may make a collection consisting of the Document and other documents
released under this License, and replace the individual copies of this
License in the various documents with a single copy that is included in
the collection, provided that you follow the rules of this License for
verbatim copying of each of the documents in all other respects.
You may extract a single document from such a collection, and distribute
it individually under this License, provided you insert a copy of this
License into the extracted document, and follow this License in all
other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS
A compilation of the Document or its derivatives with other separate
and independent documents or works, in or on a volume of a storage or
distribution medium, does not as a whole count as a Modified Version
of the Document, provided no compilation copyright is claimed for the
compilation.  Such a compilation is called an "aggregate", and this
License does not apply to the other self-contained works thus compiled
with the Document, on account of their being thus compiled, if they
are not themselves derivative works of the Document.
If the Cover Text requirement of section 3 is applicable to these
copies of the Document, then if the Document is less than one quarter
of the entire aggregate, the Document's Cover Texts may be placed on
covers that surround only the Document within the aggregate.
Otherwise they must appear on covers around the whole aggregate.

8. TRANSLATION
Translation is considered a kind of modification, so you may
distribute translations of the Document under the terms of section 4.
Replacing Invariant Sections with translations requires special
permission from their copyright holders, but you may include
translations of some or all Invariant Sections in addition to the
original versions of these Invariant Sections.  You may include a
translation of this License provided that you also include the
original English version of this License.  In case of a disagreement
between the translation and the original English version of this
License, the original English version will prevail.

9. TERMINATION

You may not copy, modify, sublicense, or distribute the Document except
as expressly provided for under this License.  Any other attempt to
copy, modify, sublicense or distribute the Document is void, and will
automatically terminate your rights under this License.  However,
parties who have received copies, or rights, from you under this
License will not have their licenses terminated so long as such
parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE
The Free Software Foundation may publish new, revised versions
of the GNU Free Documentation License from time to time.  Such new
versions will be similar in spirit to the present version, but may
differ in detail to address new problems or concerns.  See
http://www.gnu.org/copyleft/.
Each version of the License is given a distinguishing version number.
If the Document specifies that a particular numbered version of this
License "or any later version" applies to it, you have the option of
following the terms and conditions either of that specified version or
of any later version that has been published (not as a draft) by the
Free Software Foundation.  If the Document does not specify a version
number of this License, you may choose any version ever published (not
as a draft) by the Free Software Foundation.

ADDENDUM: How to use this License for your documents
To use this License in a document you have written, include a copy of
the License in the document and put the following copyright and
license notices just after the title page:
      Copyright (c)  YEAR  YOUR NAME.
      Permission is granted to copy, distribute and/or modify this document
      under the terms of the GNU Free Documentation License, Version 1.1
      or any later version published by the Free Software Foundation;
      with the Invariant Sections being LIST THEIR TITLES, with the
      Front-Cover Texts being LIST, and with the Back-Cover Texts being LIST.
      A copy of the license is included in the section entitled "GNU
      Free Documentation License".
If you have no Invariant Sections, write "with no Invariant Sections"
instead of saying which ones are invariant.  If you have no
Front-Cover Texts, write "no Front-Cover Texts" instead of
"Front-Cover Texts being LIST"; likewise for Back-Cover Texts.

If your document contains nontrivial examples of program code, we
recommend releasing these examples in parallel under your choice of
free software license, such as the GNU General Public License,
to permit their use in free software.